OpenStudy (anonymous):

Express as a single logarithm with a coefficient of 1. 7(ln(r) + ln (t))

OpenStudy (mertsj):

ln(r^7t)

OpenStudy (anonymous):

ln(r)+ln(t)=ln(rt) and 7ln(rt)=ln((rt)^7)